# Monotone Lipschitz-Gradient Denoiser: Explainability of Operator Regularization Approaches Free From Lipschitz Constant Control

### v2

@@ III-B Extension of Moreau’s Decomposition
The relation in (22) is consistent with Theorem 1 which shows $\beta$-cocoercivity of ${\rm s\mbox{-}Prox}_{\varphi}$ (see Lemma B.1 in Appendix B).
@@
We shall show that such plug-and-play methods (employing our denoiser $T=\nabla\psi$) asymptotically optimize a certain cost function in the sense of generating a vector sequence convergent to its minimizer. Here, the cost function involves an implicit “operator-inducing” regularizer which depends on $\varphi=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$. In this way, the loss function is implicitly regularized by the denoiser $T$, as opposed to the case of the traditional functional-regularization approaches which explicitly regularize the loss by some penalty functions.
@@ IV-A Forward-Backward Splitting Type Algorithm
Let $f\in{\Gamma_{0}(\mathcal{H})}$ be a $\kappa$-smooth $\rho$-strongly-convex function with $\kappa>\rho>0$. Assume that $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is a MoL-Grad denoiser (see Definition 1) for $\beta\in((\kappa-\rho)/(\kappa+\rho),1)\subsetneq(0,1)$ so that the Lipschitz constant is bounded by $\beta^{-1}<(\kappa+\rho)/(\kappa-\rho)$. Set $\mu\in[(1-\beta)/\rho,(1+\beta)/\kappa)$. Then, the following hold.
@@
    $T\circ({\rm Id}-\mu\nabla f)={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})$ (26)
    with $\beta^{-1}\in(1,2/L_{\mu\nabla\hat{f}})$, where $L_{\mu\nabla\hat{f}}:=\mu\kappa-(1-\beta)>0$ is the Lipschitz constant of $\mu\nabla\hat{f}$.
@@
    Suppose that $\mu f+\varphi$ has a minimizer in ${\mathcal{H}}$. Then, for an arbitrary $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}\subset{\mathcal{H}}$ generated by (1) converges weakly to a minimizer $\hat{x}$ of $\mu f+\varphi$. (In this case, $\varphi$ is the implicit regularizer.)
@@
$T\circ({\rm Id}-\mu\nabla f)$ $={\rm s\mbox{-}Prox}_{\varphi}\circ({\rm Id}-\mu\nabla f)$
@@
$={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f}).$
@@
2) By Theorem 2.1, (1) can be regarded as the forward-backward splitting iterate for the sum of the convex functions $\mu\hat{f}$ and $\check{\varphi}$ with the step size $\beta^{-1}\in(1,2/L_{\mu\nabla\hat{f}})$. Hence, from the standard argument of the forward-backward splitting method (see, e.g., [49, Example 17.6]), the composition ${\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})$ is an averaged nonexpansive operator with the fixed-point set ${\rm Fix}\left({{\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})}\right)=\operatornamewithlimits{argmin}_{x\in{\mathcal{H}}}[\mu\hat{f}(x)+\check{\varphi}(x)]=\operatornamewithlimits{argmin}_{x\in{\mathcal{H}}}[\mu f(x)+\varphi(x)]\neq\varnothing$, where the nonemptiness is due to the assumption. The classical Krasnosel’skĭ-Mann iterate (see Fact 1) thus ensures weak convergence of $(x_{k})_{k\in{\mathbb{N}}}$ to a fixed point which is a minimizer of $\mu f(x)+\varphi(x)$. ∎
@@ IV-B Primal-Dual Splitting Type Algorithm
We now build a framework which could be applied to a wider class of problems involving total variation, for instance, where a bounded linear operator $L:{\mathcal{H}}\rightarrow\mathcal{U}$ will be used in the algorithm. Here, ${\mathcal{H}}$ and $\mathcal{U}$ are real Hilbert spaces equipped with the inner products $\left\langle{\cdot},{\cdot}\right\rangle_{{\mathcal{H}}}$, $\left\langle{\cdot},{\cdot}\right\rangle_{\mathcal{U}}$, and the induced norms $\left\|\cdot\right\|_{{\mathcal{H}}}$, $\left\|\cdot\right\|_{\mathcal{U}}$, respectively. We define the operator norm $\left\|L\right\|:=\sup_{x\in{\mathcal{H}},\left\|x\right\|_{{\mathcal{H}}}\leq 1}\left\|Lx\right\|_{\mathcal{U}}$. The adjoint operator of $L$ is denoted by $L^{*}:\mathcal{U}\rightarrow{\mathcal{H}}$. With a nonlinear operator $T:\mathcal{U}\rightarrow\mathcal{U}$, we consider the following algorithm inspired by the Condat–Vũ algorithm [8, 9].999 An earlier study of an operator-regularization approach using the primal-dual splitting algorithm can be found in [53].
@@
is $\kappa$-smooth for some $\kappa\in{\mathbb{R}}_{++}$. Assume that $T:\mathcal{U}\rightarrow\mathcal{U}$ is a MoL-Grad denoiser for some $\psi\in\Gamma_{0}({\mathcal{H}})$ and $\beta\in(0,1)$. Set the step sizes $\sigma,\tau>0$ such that
@@
$g:=$ $~{}\Big{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big{)}\varphi=\Big{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big{)}\Big{(}\psi^{*}-\frac{1}{2}\left\|\cdot\right\|_{\mathcal{U}}^{2}\Big{)},$ (30)
$\check{g}:=$ $~{}g+\frac{\rho}{2\left\|L\right\|^{2}}\left\|\cdot\right\|_{\mathcal{U}}^{2}\in\Gamma_{0}(\mathcal{U})$ (31)
@@
The point is that the denoiser can be expressed as $T={\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ based on Theorem 1, which is further linked to the (Moreau’s) proximity operator by Proposition 3. Thanks to this link, Algorithm 1 has been shown to minimize $f+g\circ L$ with the operator-inducing function $g$ given in (30). In general, given an arbitrary operator splitting algorithm, the proximity operator(s) employed can be replaced systematically by any MoL-Grad denoiser(s) by the following steps.
@@ V-A2 Firm shrinkage
${\rm firm}_{\lambda_{1},\lambda_{2}}(x)\!:=\!\begin{cases}0&\!\!\mbox{if }\left|x\right|\leq\lambda_{1}\\ {\rm sign}(x)\frac{\lambda_{2}(\left|x\right|-\lambda_{1})}{\lambda_{2}-\lambda_{1}}&\!\!\mbox{if }\lambda_{1}<\left|x\right|\leq\lambda_{2}\\ x&\!\!\mbox{if }\left|x\right|>\lambda_{2}\end{cases}$ (37)
@@ V-A3 Garrote shrinkage
$\varphi(x):=$ $~{}\frac{1}{4}\left(\left|x\right|\sqrt{x^{2}+4\lambda^{2}}-x^{2}\right)$
@@ V-B Shrinkage Operators for Vector
[…]
@@
[…]
@@
Let $\nabla\psi:{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ be a firmly-nonexpansive activation operator (see Section II-C). In fact, many of the known activation functions are firmly nonexpansive [2], such as ReLU, sigmoid, and softmax. For instance, ReLU ${\mathbb{R}}\rightarrow{\mathbb{R}}:x\mapsto\begin{cases}x&\mbox{ if }x\in{\mathbb{R}}_{+}\\ 0&\mbox{ otherwise}\end{cases}$ can be expressed as the proximity operator ${\rm s\mbox{-}Prox}_{\varphi}$ of the indicator function $\varphi:=\iota_{{\mathbb{R}}_{+}}:{\mathbb{R}}\rightarrow\{0,+\infty\}:x\mapsto\begin{cases}0&\mbox{ if }x\in{\mathbb{R}}_{+}\\ +\infty&\mbox{ otherwise,}\end{cases}$ where ${\rm s\mbox{-}Prox}_{\varphi}$ is firmly nonexpansive owing to the convexity of $\iota_{{\mathbb{R}}_{+}}$ (see Section II-E). Let ${\boldsymbol{W}}\in{\mathbb{R}}^{m\times n}$ be a nonzero weight matrix. Then, a weight-tied neural network [56] $T:={\boldsymbol{W}}^{\top}\circ\nabla\psi\circ{\boldsymbol{W}}=\nabla(\psi\circ{\boldsymbol{W}})$ can be expressed as $T={\rm s\mbox{-}Prox}_{\varphi}$ with $\varphi:=(\psi\circ{\boldsymbol{W}})^{*}-(1/2)\left\|\cdot\right\|_{2}^{2}$. Here, if $\kappa:=\|{\boldsymbol{W}}^{\top}{\boldsymbol{W}}\|>1$, $T$ is a MoL-Grad denoiser, and $\varphi$ is $(1-\kappa^{-1})$-weakly convex. If $\kappa\in(0,1]$, in contrast, $T$ is Moreau’s proximity operator of the convex function $\varphi$. More practical network architectures are discussed in [57], where multi-layered neural networks with tied nonnegative weights involving residual skip connections are shown to be MoL-Grad denoisers.
Example 4 (Debiased OSCAR)
@@
$T_{\lambda_{1},\lambda_{2},\omega,\eta}^{{\rm DOSCAR}}:=\omega{\rm s\mbox{-}Prox}_{\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}-(1-\omega){\rm s\mbox{-}Prox}_{\eta\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}$ (44)
@@ V-C Prior Art I: Plug-and-Play Methods
TABLE I: Comparisons of the existing denoisers for the plug-and-play method in terms of (i) the condition on the Lipschitz constant L()($\cdot$)( ⋅ ), (ii) smoothness of the implicit regularizer $\varphi$, and (iii) the range of the weak-convexity constant of $\varphi$
@@
Gradient-step (GS) denoiser: The GS denoiser has been studied in the literature [68, 69, 37]. Among those studies, the most relevant results would be the one in [37, Proposition 3.1], where ${\rm Id}-\nabla\psi$ is assumed to be contraction ($\kappa$-Lipschitz continuous for $\kappa<1$).101010 The arguments of [37, Proposition 3.1] rely essentially on [70]. Because of this restrictive assumption, the result in [37, Proposition 3.1] has several limitations: (i) $T$ possesses a positive definite Jacobian matrix at every point, and (ii) $\varphi$ is smooth and $\eta_{\rm GS}$-weakly convex for $\eta_{\rm GS}\in(0,1/2)$ ($\psi$ is strongly convex). In particular, the smoothness of $\varphi$ could be a strict limitation (see Remark 5 below). The implicit regularizer given in [37, Proposition 3.1] can be reproduced as a special case of Theorem 1 straightforwardly.
@@
Regularization by denoising (RED): The RED approach [67] relates specific algorithms to a variational optimization problem. We emphasize that it is not such an approach that relates a denoiser itself directly to an implicit regularizer. In this approach, the denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is assumed to possess (i) nonexpansiveness, (ii) symmetric Jacobian $JT({\boldsymbol{x}})$ at every point ${\boldsymbol{x}}\in{\mathcal{H}}$, and (iii) the so-called local homogeneity: $T(c{\boldsymbol{x}})=cT({\boldsymbol{x}})$ for every $c\in[1-\epsilon,1+\epsilon]$ for a sufficiently small $\epsilon>0$. Here, (iii) is a key assumption to ensure $T({\boldsymbol{x}})=[JT({\boldsymbol{x}})]{\boldsymbol{x}}$, from which it can be shown that $\nabla h_{\rm RED}({\boldsymbol{x}})={\boldsymbol{x}}-T({\boldsymbol{x}})$ for $h_{\rm RED}({\boldsymbol{x}}):=(1/2){\boldsymbol{x}}^{\top}({\boldsymbol{x}}-T({\boldsymbol{x}}))$. We stress here that the relation between the denoiser $T$ and the regularizer $h_{\rm RED}$ is essentially different from our case. It has been pointed out that the RED algorithm does not actually minimize the variational objective [72], and another approach called consensus equilibrium has also been proposed as an optimization-free generalization [73].
Characterization of the limit point for plug-and-play method: In [35, Proposition 3.1], a characterization of the limit point of the vector sequence generated by the forward-backward splitting algorithm using the plug-and-play method has been presented through variational inequalities for a general family of maximally monotone operator. While the scope therein is wider in the sense that their analysis covers a general case rather than focusing solely on optimization, the operator considered therein is neural network denoisers. In addition, we explicitly showed in Section IV-B that, when the plug-and-play is applied to the primal-dual splitting algorithm, it is still possible to ensure the convergence to an optimal point with a slight modification. We repeat that the operator is not necessarily nonexpansive in our case, whereas it is supposed to be firmly nonexpansive in [35, Proposition 3.1].
@@
Table I shows comparisons of the proposed and existing approaches. To the best of the authors’ knowledge, MoL-Grad is the first denoiser that requires no control of its Lipschitz constant to guarantee the convergence of the operator-regularization (PnP) algorithms. In addition, the implicit regularizer $\varphi(=\psi^{*}-(1/2)\left\|\cdot\right\|^{2})$ induced by MoL-Grad denoisers is not limited to smooth ones. This means that the class of MoL-Grad denoiser is sufficiently large to contain those denoisers inducing nonsmooth regularizers. Note here that, for $\varsigma\in{\mathbb{R}}_{++}$, (i) $\varphi$ is $\varsigma$-smooth $\Rightarrow$ (ii) $\varphi+(1/2)\left\|\cdot\right\|^{2}(=\psi^{*})$ is $(\varsigma+1)$-smooth $\Leftrightarrow$ (iii) $\psi$ is $1/(\varsigma+1)$-strongly convex $\Rightarrow$ (iv) $T=\nabla\psi$ is $1/(\varsigma+1)$-strongly monotone (see Section II-B for the equivalence (ii) $\Leftrightarrow$ (iii)). This implies that the MoL-Grad denoiser $T$ needs not to be strongly monotone. Nonsmoothness of regularizers has actually played a crucial role in sparse modeling. For instance, the firm/garrote shrinkage presented in Section V-A induces the nonsmooth regularizer given in (38) or (40). For the weight-tied neural networks discussed in Example 3, moreover, adoption of a saturating activation function in the network such as ReLU makes the associated regularizer nonsmooth, because $\nabla^{2}\psi({\boldsymbol{y}})=JT({\boldsymbol{y}})={\boldsymbol{O}}$ at some ${\boldsymbol{y}}\in{\mathbb{R}}^{m}$.

### v3

@@ III-B Extension of Moreau’s Decomposition
The relation in (22) implies $\beta$-cocoercivity of ${\rm s\mbox{-}Prox}_{\varphi}$ (see Lemma B.1 in Appendix B), which is consistent with Theorem 1.
@@
We shall show that such plug-and-play methods (employing our denoiser $T=\nabla\psi$) asymptotically optimize a certain cost function in the sense of generating a vector sequence convergent to its minimizer. Here, the cost function involves an implicit “operator-induced” regularizer which depends on $\varphi=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$. In this way, the loss function is implicitly regularized by the denoiser $T$, as opposed to the case of the traditional functional-regularization approaches which explicitly regularize the loss by some penalty functions.
@@ IV-A Forward-Backward Splitting Type Algorithm
Let $f\in{\Gamma_{0}(\mathcal{H})}$ be a $\kappa$-smooth $\rho$-strongly-convex function with $\kappa>\rho>0$. Assume that $T:=\nabla\psi:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is a MoL-Grad denoiser (see Definition 1) for $\beta\in((\kappa-\rho)/(\kappa+\rho),1)\subsetneq(0,1)$ so that the Lipschitz constant is bounded by $\beta^{-1}<(\kappa+\rho)/(\kappa-\rho)$. Set $\mu\in[(1-\beta)/\rho,(1+\beta)/\kappa)$. Let $\varphi:=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$. Then, the following hold.
@@
[…]
@@
    Suppose that $\mu f+\varphi$ has a minimizer in $\mathcal{H}$}. Then, for an arbitrary $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}\subset{\mathcal{H}}$ generated by (1) converges weakly to a minimizer $\hat{x}\in{\mathcal{H}}$ of $\mu f+\varphi$. (In this case, $\varphi$ is the implicit regularizer.)
@@
$\hskip-10.00002ptT\circ({\rm Id}-\mu\nabla f)$ $={\rm s\mbox{-}Prox}_{\varphi}\circ({\rm Id}-\mu\nabla f)$
@@
[…]
@@
[…]
@@ IV-B Primal-Dual Splitting Type Algorithm
We now build a framework which could be applied to a wider class of problems involving total variation, for instance, where a bounded linear operator $L:{\mathcal{H}}\rightarrow\mathcal{U}$ will be used in the algorithm. Here, $\mathcal{H}$} and $\mathcal{U}$ are real Hilbert spaces equipped with the inner products $\left\langle{\cdot},{\cdot}\right\rangle_{{\mathcal{H}}}$, $\left\langle{\cdot},{\cdot}\right\rangle_{\mathcal{U}}$, and the induced norms $\left\|\cdot\right\|_{{\mathcal{H}}}$, $\left\|\cdot\right\|_{\mathcal{U}}$, respectively. We define the operator norm $\left\|L\right\|:=\sup_{x\in{\mathcal{H}},\left\|x\right\|_{{\mathcal{H}}}\leq 1}\left\|Lx\right\|_{\mathcal{U}}$. The adjoint operator of $L$ is denoted by $L^{*}:\mathcal{U}\rightarrow{\mathcal{H}}$. With a nonlinear operator $T:\mathcal{U}\rightarrow\mathcal{U}$, we consider the following algorithm inspired by the Condat–Vũ algorithm [8, 9]. [^9] An earlier study of an operator-regularization approach using the primal-dual splitting algorithm can be found in [53].
@@
is $\kappa$-smooth for some $\kappa\in{\mathbb{R}}_{++}$. Assume that $T:=\nabla\psi:\mathcal{U}\rightarrow\mathcal{U}$ is a MoL-Grad denoiser for some $\psi\in\Gamma_{0}({\mathcal{H}})$ and $\beta\in(0,1)$. Set the step sizes $\sigma,\tau>0$ such that
@@
$\hskip-20.00003ptg:=$ $~\Big(\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big)\varphi=\Big(\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big)\Big(\psi^{*}-\frac{1}{2}\left\|\cdot\right\|_{\mathcal{U}}^{2}\Big),$ (30)
$\hskip-20.00003pt\check{g}:=$ $~g+\frac{\rho}{2\left\|L\right\|^{2}}\left\|\cdot\right\|_{\mathcal{U}}^{2}\in\Gamma_{0}(\mathcal{U})$ (31)
@@
The point is that the denoiser can be expressed as $T={\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ based on Theorem 1, which is further linked to the (Moreau’s) proximity operator by Proposition 3. Thanks to this link, Algorithm 1 has been shown to minimize $f+g\circ L$ with the operator-induced function $g$ given in (30). In general, given an arbitrary operator splitting algorithm, the proximity operator(s) employed can be replaced systematically by any MoL-Grad denoiser(s) by the following steps.
@@ V-A2 Firm shrinkage
$\hskip-1.00006pt{\rm firm}_{\lambda_{1},\lambda_{2}}(x)\!:=\!\begin{cases}0&\!\!\mbox{if }\left|x\right|\leq\lambda_{1}\\ {\rm sign}(x)\frac{\lambda_{2}(\left|x\right|-\lambda_{1})}{\lambda_{2}-\lambda_{1}}&\!\!\mbox{if }\lambda_{1}<\left|x\right|\leq\lambda_{2}\\ x&\!\!\mbox{if }\left|x\right|>\lambda_{2}\end{cases}$ (37)
@@ V-A3 Garrote shrinkage
$\hskip-10.00002pt\varphi(x):=$ $~\frac{1}{4}\left(\left|x\right|\sqrt{x^{2}+4\lambda^{2}}-x^{2}\right)$
@@ V-B Shrinkage Operators for Vector
Using this notion, the MC function $\varphi_{\lambda_{2}}^{\rm MC}$ can be expressed as the Moreau enhanced model of the absolute-value function $\left|\cdot\right|$ of index $\lambda_{2}$; i.e., $\varphi_{\lambda_{2}}^{\rm MC}(x)=(\left|\cdot\right|)_{\lambda_{2}}$ [29, Proposition 12].
@@
Let $\mathcal{H}$} be an arbitrary Hilbert space equipped with a norm $\left\|\cdot\right\|$. For positive constants $\lambda_{1}\in{\mathbb{R}}_{++}$ and $\lambda_{2}\in(\lambda_{1},+\infty)$, we define the vector firm-shrinkage operator:
@@
Let $\nabla\psi:{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ be a firmly-nonexpansive activation operator (see Section II-C). In fact, many of the known activation functions are firmly nonexpansive [2], such as ReLU, sigmoid, and softmax. For instance, ReLU ${\mathbb{R}}\rightarrow{\mathbb{R}}:x\mapsto\begin{cases}x&\mbox{ if }x\in{\mathbb{R}}_{+}\\ 0&\mbox{ otherwise}\end{cases}$ can be expressed as the proximity operator ${\rm s\mbox{-}Prox}_{\varphi}$ of the indicator function $\varphi:=\iota_{{\mathbb{R}}_{+}}:{\mathbb{R}}\rightarrow\{0,+\infty\}:x\mapsto\begin{cases}0&\mbox{ if }x\in{\mathbb{R}}_{+}\\ +\infty&\mbox{ otherwise,}\end{cases}$ where ${\rm s\mbox{-}Prox}_{\varphi}$ is firmly nonexpansive owing to the convexity of $\iota_{{\mathbb{R}}_{+}}$ (see Section II-E). Let ${\boldsymbol{W}}\in{\mathbb{R}}^{m\times n}$ be a nonzero weight matrix. Then, a weight-tied neural network [56] $T:={\boldsymbol{W}}^{\top}\circ\nabla\psi\circ{\boldsymbol{W}}=\nabla(\psi\circ{\boldsymbol{W}})$ can be expressed as $T={\rm s\mbox{-}Prox}_{\varphi}$ with $\varphi:=(\psi\circ{\boldsymbol{W}})^{*}-(1/2)\left\|\cdot\right\|_{2}^{2}$. Here, if $\kappa:=\|{\boldsymbol{W}}^{\top}{\boldsymbol{W}}\|>1$, $T$ is a MoL-Grad denoiser, and $\varphi$ is $(1-\kappa^{-1})$-weakly convex. If $\kappa\in(0,1]$, in contrast, $T$ is Moreau’s proximity operator of the convex function $\varphi$. More practical network architectures are discussed in [57], where multi-layered neural networks with tied nonnegative weights are shown to be MoL-Grad denoisers.
Example 4 (Debiased OSCAR [51])
@@
$T_{\lambda_{1},\lambda_{2},\omega,\eta}^{{\rm DOSCAR}}:=\omega{\rm s\mbox{-}Prox}_{\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}-(\omega-1){\rm s\mbox{-}Prox}_{\eta\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}$ (44)
@@ V-C Prior Art I: Plug-and-Play Methods
TABLE I: Comparisons of the existing denoisers for the plug-and-play method in terms of (i) the condition on the Lipschitz constant L()($\cdot$), (ii) smoothness of the implicit regularizer $\varphi$, and (iii) the range of the weak-convexity constant of $\varphi$
@@
Gradient-step (GS) denoiser: The GS denoiser has been studied in the literature [68, 69, 37]. Among those studies, the most relevant results would be the one in [37, Proposition 3.1], where $T:={\rm Id}-\nabla g$ is assumed to be contraction ($\kappa$-Lipschitz continuous for $\kappa<1$). [^10] The arguments of [37, Proposition 3.1] rely essentially on [70]. Because of this restrictive assumption, the result in [37, Proposition 3.1] has several limitations: (i) $T$ possesses a positive definite Jacobian matrix at every point, and (ii) $\varphi$ is smooth and $\eta_{\rm GS}$-weakly convex for $\eta_{\rm GS}\in(0,1/2)$ ($\psi:=(1/2)\left\|\cdot\right\|^{2}-g$ is strongly convex and smooth at the same time). In particular, the smoothness of $\varphi$ could be a strict limitation (see Remark 5 below). The implicit regularizer given in [37, Proposition 3.1] can be reproduced as a special case of Theorem 1 straightforwardly.
@@
Regularization by denoising (RED): The RED approach [67] relates specific algorithms to a variational optimization problem. We emphasize that it is not such an approach that relates a denoiser itself directly to an implicit regularizer. In this approach, the denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is assumed to possess (i) nonexpansiveness, (ii) symmetric Jacobian $JT({\boldsymbol{x}})$ at every point ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, and (iii) the so-called local homogeneity: $T(c{\boldsymbol{x}})=cT({\boldsymbol{x}})$ for every $c\in[1-\epsilon,1+\epsilon]$ for a sufficiently small $\epsilon>0$. Here, (iii) is a key assumption to ensure $T({\boldsymbol{x}})=[JT({\boldsymbol{x}})]{\boldsymbol{x}}$, from which it can be shown that $\nabla h_{\rm RED}({\boldsymbol{x}})={\boldsymbol{x}}-T({\boldsymbol{x}})$ for $h_{\rm RED}({\boldsymbol{x}}):=(1/2){\boldsymbol{x}}^{\top}({\boldsymbol{x}}-T({\boldsymbol{x}}))$. We stress here that the relation between the denoiser $T$ and the regularizer $h_{\rm RED}$ is essentially different from our case. It has been pointed out that the RED algorithm does not actually minimize the variational objective [72], and another approach called consensus equilibrium has also been proposed as an optimization-free generalization [73].
Characterization of the limit point for plug-and-play method: In [35, Proposition 3.1], a characterization of the limit point of the vector sequence generated by the forward-backward splitting algorithm using the plug-and-play method has been presented through variational inequalities for a general family of maximally monotone operator. Their analysis covers a general case rather than focusing solely on optimization. However, the operator is supposed to be firmly nonexpansive in [35, Proposition 3.1], whereas it is not necessarily nonexpansive in our case. In addition, we explicitly showed in Section IV-B that, when the plug-and-play is applied to the primal-dual splitting algorithm, it is still possible to ensure the convergence to an optimal point with a slight modification.
@@
Table I shows comparisons of the proposed and existing approaches. To the best of the authors’ knowledge, MoL-Grad is the first denoiser that requires no control of its Lipschitz constant to guarantee the convergence of the operator-regularization (PnP) algorithms. In addition, the implicit regularizer $\varphi(=\psi^{*}-(1/2)\left\|\cdot\right\|^{2})$ induced by MoL-Grad denoisers is not limited to smooth ones. This means that the class of MoL-Grad denoiser is sufficiently large to contain those denoisers inducing nonsmooth regularizers.
Nonsmoothness of regularizers has actually played a crucial role in sparse modeling. For instance, the firm/garrote shrinkage presented in Section V-A induces the nonsmooth regularizer given in (38) or (40). For the weight-tied neural networks discussed in Example 3, moreover, adoption of a saturating activation function in the network such as ReLU makes the associated regularizer nonsmooth, because $\nabla^{2}\psi({\boldsymbol{y}})=JT({\boldsymbol{y}})={\boldsymbol{O}}$ at some ${\boldsymbol{y}}\in{\mathbb{R}}^{m}$.
We finally mention that, for $\varsigma\in{\mathbb{R}}_{++}$, (i) $\varphi$ is $\varsigma$-smooth $\Rightarrow$ (ii) $\varphi+(1/2)\left\|\cdot\right\|^{2}(=\psi^{*})$ is $(\varsigma+1)$-smooth $\Leftrightarrow$ (iii) $\psi$ is $1/(\varsigma+1)$-strongly convex $\Rightarrow$ (iv) $T=\nabla\psi$ is $1/(\varsigma+1)$-strongly monotone (see Section II-B for the equivalence (ii) $\Leftrightarrow$ (iii)). This implies that the GS denoiser is restricted to strongly monotone operators, while the MoL-Grad denoiser is free from such a restriction.